# Sensing for Free: Learn to Localize More Sources than Antennas without Pilots

### v1

@@ V-A3 Training Settings
We train the neural networks by using stochastic gradient descent (SGD) with a batch size of 4096. The models are trained for 100 epochs with the one-cycle scheduler of learning rates. For the benchmarks, we use the same learning rates as suggested in [31] (0.05 for DCR-T, 0.01 for DCR-G-Fro, 0.005 for DCR-G-Aff, and 0.1 for SRL-GD). For the proposed Snap-TF algorithm, we utilize a learning rate of 0.001. We generate $2\times 10^{6}$ training, $6\times 10^{4}$ validation, and $1\times 10^{4}$ testing samples, respectively. We use PyTorch to train and test all the models on one Nvidia A40 GPU.
@@ V-B Performance
In Fig. 3, we plot the MSE of DOA estimation versus SNR for $M=5$ and $N=9$ under both Gaussian (the top panel) and non-Gaussian modulated sources (the bottom panel) with the number of sources $K=3,9$ arranged from left to right. These respectively represent cases where the number of sources is smaller and larger than the number of antennas. The model is trained over a mixture of SNR levels drawn uniformly from $[-20,20]$ dB, while tested in an SNR range of $[-30,20]$ dB.
@@
Third, increasing the number of sources from $K=3$ to $K=9$ raises the MSE for all methods, since $K=3$ (smaller than $M-1$) falls within the physical array’s degree of freedom while $K=9$ (larger than $M-1$) require exploiting the virtual array to resolve the sources. In both cases, Snap-TF consistently delivers better DOA estimation accuracy than all the benchmark methods. The extended localization capability is particularly useful for dense 6G networks with an ever-increasing number of sources that can greatly exceed the number of physical antennas, such as low-altitude UAV swarms which have immediate commercial values and are mentioned in discussions for 3GPP release 20 [6].
@@ V-C Generalization Capability
In Fig. 4, we show the generalization capability of the proposed Snap-TF algorithms to practical mismatches, including the number of snapshots $T$ and the minimum angle separation $\Delta_{\bm{\theta},\min}$. All the methods shown in the figure are trained under $T=50$ and $\Delta_{\bm{\theta},\min}=\frac{\pi}{60}$, and directly tested in different cases. During testing, the received SNR is 10 dB, and the sources transmit 16QAM signals. As shown in the figure, the Snap-TF algorithm remains remarkably robust and consistently outperforms the benchmarks in terms of generalization performance.
Figure 4: Generalization performance to different numbers of snapshots $T$ and minimum angle separations $\Delta_{\bm{\theta},\min}$. The model is trained under $T=50$ and $\Delta_{\bm{\theta},\min}=\frac{\pi}{60}$, and directly tested in all the different settings.
In the top panel, we evaluate the generalization performance across different snapshot counts $T$ ranging from 10 to 100, maintaining a fixed minimum angle separation of $\Delta_{\bm{\theta},\min}=\frac{\pi}{60}$. Our Snap-TF algorithm demonstrates exceptional robustness across this range. Despite training at $T=50$, Snap-TF exhibits remarkable generalization when operating with fewer snapshots ($T<50$) and continues to achieve superior performance gains as snapshot count increases ($T>50$), consistently outperforming all benchmark methods across the entire range. This outstanding generalization capability highlights a key strength of Snap-TF architecture, i.e., their inherent ability to adapt to varying snapshot counts while maintaining robust performance.
In the bottom panel, we evaluate generalization performance across varying minimum angle separations, testing $\Delta_{\bm{\theta},\min}$ values from $\frac{\pi}{180}$ (challenging closely spaced sources) to $\frac{6\pi}{180}$ (separated sources) at a fixed $T=50$. As expected, the DOA estimation task becomes increasingly difficult as $\Delta_{\bm{\theta},\min}$ decreases, presenting a natural challenge for all methods. Our Snap-TF algorithm excels across the entire range of $\Delta_{\bm{\theta},\min}$, consistently delivering the lowest MSE, including the challenging closely spaced scenarios.
@@ V-E Complexity
As seen in Fig. 6(a), the proposed Snap-TF algorithm consistently achieves sub-millisecond level runtime under both system configurations, approximately 1.103 ms for the small-scale case ($M=5$, $N=9$) and 1.112 ms for the large-scale case ($M=15$, $N=79$). By contrast, the benchmark DL-based subspace methods, which rely on matrix decompositions, exhibit a $\times2$ slowdown over Snap-TF in the small-scale case but suffer a dramatic 31.3×31.3$\times$31.3 × increase in runtime when scaled up. This disparity underscores the heavy cost of the benchmarks, including $O(N^{3})$ operations such as matrix decomposition and inversion, making them computationally prohibitive in large-scale MIMO systems. By contrast, Snap-TF circumvents this bottleneck by eliminating the need for explicit matrix decomposition. It maintains a moderate runtime even in large-scale setups, and is suitable for real-time applications.
In Fig. 6(b), we compare the number of trainable parameters. The number for the benchmark is around 11 million and 17 million in the small-scale and large-scale setups, respectively, while that of the proposed Snap-TF algorithm (with 3 layers) is around 0.356 million in both setups. The proposed algorithm is 31.1×31.1$\times$31.1 × and 49.7×49.7$\times$49.7 × more efficient in parameters, respectively, which is lightweight for practical deployment.
@@ V-F Ablation Studies
We conduct ablation studies to demonstrate Snap-TF’s capability to effectively learn and exploit higher-order statistical for DOA estimation. In the case of Gaussian signals, $k=2$ (second) order statistics are the sufficient statistics. As discussed in Section IV, only $L=\text{$\lceil\log_{2}(k+1)\rceil$}=2$ single-head transformer layers are sufficient. This is verified by the simulation results in Fig. 7(a), where we observe the performance starts to saturate when the model uses more than two layers. The saturation behavior provides evidence that Snap-TF has successfully captured all the relevant statistical information available in the Gaussian case, with additional layers offering diminishing returns. Similarly, in the non-Gaussian case, 16QAM symbols have non-zero $k=4$ (fourth) order cumulants [33], and hence $L=\text{$\lceil\log_{2}(k+1)\rceil$}=3$ layers are required. The results in Fig. 7(b) verify our hypothesis, which show that the performance keeps improving in the first three layers, and saturates thereafter. The performance pattern matches our prediction, and confirms that Snap-TF effectively captures the higher-order statistics required for non-Gaussian DOA estimation. This also suggests a lightweight 3-layer model is sufficient for pilot-free ISAC in cellular systems based on unknown uplink data symbols. We also label the number of parameters in the unit of million in the legends.
@@ VI Further Application: Sensing-Assisted MU-MIMO Beam Management
Beam sweeping is an important step of beam management to establish initial alignment. In the downlink, the BS will transmit synchronization signal blocks sequentially across various spatial directions, effectively “sweeping” its coverage range. The users then feedback the index of the strongest beams they receive to the BS to facilitate beam alignment. In the beam sweeping stage, the BS needs to sweep its whole coverage range based on a pre-defined codebook, for which a widely used choice is the discrete Fourier transform (DFT) codebook. Advanced codebook design has also been extensively studied to reduce the beam sweeping overhead and improve the performance.
@@ VI-B Performance
Fig. 9 illustrates the system throughput as a function of the channel coherence time $T_{\text{c}}$ for three beam sweeping strategies. In Fig. 9 (a) and (b), DOA estimation from unknown uplink data symbols is performed at SNRs of -5 dB and 5 dB, respectively. Each data point in the plot is averaged over 10,000 Monte Carlo trials. The proposed codebook + sensing (pruned sweeping) approach consistently outperforms the conventional codebook-only method across all coherence time lengths, with the most pronounced advantage in the short coherence regime. This gain is attributed to the significant reduction of beam sweeping candidates enabled by the sensed uplink DOAs, which effectively lowers training overhead. The beam sweeping overhead for full sweeping is $T_{\text{train}}=64$, while the overhead of sensing pruned sweeping is significantly smaller, only leading to an average sweeping overhead of $T_{\text{train}}=5.0$ in (a) and $T_{\text{train}}=2.9$ in (b), which are respectively 12.8×12.8$\times$12.8 × and 22.1×22.1$\times$22.1 × times fewer compared to full sweeping. As the coherence interval $T_{\text{c}}$ increases, the performance of the full sweeping baseline gradually converges toward that of the proposed sensing-aided pruned sweeping scheme.
@@
VII Conclusion
This paper demonstrates the novel paradigm of “sensing for free”—enabling multi-source localization at the BS based on unknown uplink data symbols, without dedicated pilots or additional overhead. Sparse arrays prove particularly promising for this paradigm, as they have the potential to localize more sources than antennas by exploiting their enlarged virtual array. Our attention-only transformer directly processes raw signal snapshots to exploit higher-order statistics for grid-less end-to-end multi-source DOA estimation, circumventing the computational bottlenecks and performance degradation inherent in existing DL-based subspace methods. These accurate DOA estimates can enable significant pruning of downlink beam sweeping candidates, and hence substantially reduce beam training overhead and enhance system throughput at no extra cost. By leaving frame structures and waveforms untouched, the paradigm integrates seamlessly into existing wireless systems and provides inspiration for next-generation 3GPP standardization efforts. Overall, this work reveals how uplink data symbols can be harnessed for free to unlock the substantial sensing capabilities, motivating future research into the untapped potential of regular data symbols for ISAC.

### v2

@@ V-A3 Training Settings
We train the neural networks by using stochastic gradient descent (SGD) with a batch size of 4096. The models are trained for 100 epochs with the one-cycle scheduler of learning rates. For the benchmarks, we use the same learning rates as suggested in [34] (0.05 for DCR-T, 0.01 for DCR-G-Fro, 0.005 for DCR-G-Aff, and 0.1 for SRL-GD). For the proposed Snap-TF algorithm, we utilize a learning rate of 0.001. For each $K\in\{1,2,\ldots,K_{\text{max}}\}$, we generate $2\times 10^{6}$ training, $6\times 10^{4}$ validation, and $1\times 10^{4}$ testing samples, respectively. For $K_{\text{max}}=9$, this leads to a total size of $9\times 2\times 10^{6}$, $9\times 6\times 10^{4}$, and $9\times 1\times 10^{4}$ samples for training, validation, and testing, respectively. We use PyTorch to train and test all the models on one Nvidia A40 GPU.
@@ V-B Performance
In Fig. 5, we plot the MSE of DOA estimation versus SNR for $M=5$ and $N=9$ under both Gaussian (the top panel) and non-Gaussian modulated sources (the bottom panel) with the number of sources $K=3,9$ arranged from left to right. These respectively represent cases where the number of sources is smaller and larger than the number of antennas. The model is trained over a mixture of SNR levels drawn uniformly from [20,20][-20,20] dB, while tested in an SNR range of [30,20][-30,20] dB.
@@
Third, increasing the number of sources from $K=3$ to $K=9$ raises the MSE for all methods, since $K=3$ (smaller than M1M-1) falls within the physical array’s degree of freedom while $K=9$ (larger than M1M-1) require exploiting the virtual array to resolve the sources. In both cases, Snap-TF consistently delivers better DOA estimation accuracy than all the benchmark methods. This extended localization capability is particularly useful for dense 6G networks with more sources than antennas (e.g., low-altitude UAV swarms) [7].
@@ V-C Generalization Capability
In Fig. 6, we show the generalization capability of the proposed Snap-TF algorithms to practical mismatches, including the number of snapshots $T$ and the different modulation types. All the methods shown in the figure are trained under $T=50$ and $\Delta_{\bm{\theta},\min}=\frac{\pi}{60}$, and directly tested in different cases without retraining or tuning. During testing, the SNR is set as 10 dB.
Figure 6: Generalization performance to different numbers of snapshots $T$ and different modulation types. The model is trained under $T=50$ and $\Delta_{\bm{\theta},\min}=\frac{\pi}{60}$, and directly tested in all the different settings.
In the top panel, we evaluate generalization across the numbers of snapshots $T$ from 10 to 100 while fixing the minimum angle separation at $\Delta_{\bm{\theta},\min}=\frac{\pi}{60}$. Despite being trained at $T=50$, Snap-TF achieves superior performance for $20\leq T\leq 100$, demonstrating adaptability to varying numbers of snapshots while maintaining robustness. A notable exception is $T=10$ in Fig. 6(a), where Snap-TF slightly underperforms the best benchmark. We conjecture that, because $T=10$ is very small, the sample estimates of higher-order statistics exploited by Snap-TF have high variance. Moreover, this test uses $K<M$ (fewer sources than antennas), a regime in which second-order statistics (used by the benchmarks) already perform well. As a result, higher-order information could be misleading at such small $T$, leading to a disadvantage for Snap-TF at $T=10$. The slight fluctuations are attributed to the fact that these are out-of-distribution cases that the models were not trained on.
In the bottom panel, we evaluate the generalization performance of Snap-TF to different, or even mixed, modulation types that it was not trained on. Specifically, we set up a “mixed modulation type” case where each source transmits in either QPSK or 16QAM modulation. We draw a Bernoulli distributed mask, i.e., $\mathrm{Bernoulli}(0.5)$, for every symbol transmitted by each source in each snapshot. If the random mask is 1, the source transmits a QPSK symbol. Otherwise, if the mask equals 0, the source transmits a 16QAM symbol instead. We test three variants of the proposed Snap-TF algorithm on such a “mixed modulation type” case. The left-hand-side of the arrow denotes the dataset where the model is trained on, while the right-hand-side denotes the testing dataset of the model. For example, 16QAM $\rightarrow$ QPSK+16QAM denotes the case where a model trained on 16QAM sources, but tested on mixed modulation types of QPSK+16QAM sources. Simulation results show that under both $K=3$ and $K=9$ sources, the curve trained on 16QAM only is very close to the model trained on the mixed QPSK+16QAM data. This suggests that after trained on a single non-Gaussian, discrete constellation with amplitude variation, Snap-TF has the potential to generalize to different, and even mixed, modulation types. In contrast, the Gaussian $\rightarrow$ QPSK+16QAM model shows a notable performance degradation at medium to high SNR, which is likely attributed to the absence of higher-order statistics when the model is trained on Gaussian signals.
@@ V-E Complexity
As seen in Fig. 8(a), the proposed Snap-TF algorithm consistently achieves sub-millisecond level runtime under both system configurations, approximately 1.103 ms for the small-scale case ($M=5$, $N=9$) and 1.112 ms for the large-scale case ($M=15$, $N=79$). By contrast, the benchmark DL-based subspace methods, which rely on matrix decompositions, exhibit a $2.2\times$ slowdown over Snap-TF in the small-scale case but suffer a dramatic $31.3\times$ increase in runtime when scaled up. This disparity underscores the heavy cost of the benchmarks, including $O(N^{3})$ operations such as matrix decomposition and inversion, making them computationally prohibitive in large-scale MIMO systems. By contrast, Snap-TF circumvents this bottleneck by eliminating the need for explicit matrix decomposition. It maintains a moderate runtime even in large-scale setups, and is suitable for real-time applications.
In Fig. 8(b), we compare the number of trainable parameters. The number for the benchmark is around 11 million and 17 million in the small-scale and large-scale setups, respectively, while that of the proposed Snap-TF algorithm (with 3 layers) is around 0.356 million in both setups. The proposed algorithm is $31.1\times$ and $49.7\times$ more efficient in parameters, respectively, which is lightweight for practical deployment.
@@ V-F Ablation Studies
We first conduct an ablation study to identify the sources of performance gains by isolating two key factors: (i) raw snapshot processing versus sample SCM-based input, and (ii) end-to-end (E2E) training versus two-stage algorithm (learning covariance reconstruction followed by Root-MUSIC). We compare five methods in Fig. 9: the proposed Snap-TF (raw snapshots + E2E), SRL-GD [34] (sample SCM + two-stage), DCR-E2E (sample SCM + E2E), an Oracle bound [^7] This is obtained by applying Root-MUSIC to the sample SCM of a physical 10-element ULA (matching the virtual aperture of our 5-element MRA) under identical conditions ($\text{SNR}=20$ dB, $T=50$). This represents an ideal case and may be loose as it uses physical 10-element ULA., and the a priori bound (APB) [64] (performance bound at low SNRs).
Figure 9: Ablation studies of the reasons behind the performance gains.
[…]
We also conduct ablation studies to demonstrate Snap-TF’s capability to effectively learn and exploit higher-order statistical for DOA estimation. In the case of Gaussian signals, $k=2$ (second) order statistics are the sufficient statistics. As discussed in Section IV, only $L=\text{$\lceil\log_{2}(k+1)\rceil$}=2$ single-head transformer layers are sufficient. This is verified by the simulation results in Fig. 10(a), where we observe the performance starts to saturate when the model uses more than two layers. The saturation behavior provides evidence that Snap-TF has successfully captured all the relevant statistical information available in the Gaussian case, with additional layers offering diminishing returns. Similarly, in the non-Gaussian case, 16QAM symbols have non-zero $k=4$ (fourth) order cumulants [36], and hence $L=\text{$\lceil\log_{2}(k+1)\rceil$}=3$ layers are required. The results in Fig. 10(b) verify our hypothesis, which show that the performance keeps improving in the first three layers, and saturates thereafter. The performance pattern matches our prediction, and confirms that Snap-TF effectively captures the higher-order statistics required for non-Gaussian DOA estimation. This also suggests a lightweight 3-layer model is sufficient for multi-source localization in cellular systems based on unknown uplink data symbols. We also label the number of parameters in the unit of million in the legends.
@@
V-G Multipath Propagation and Coherent Sources
We now discuss the impact of multipath propagation. As an example, assume two signals are coherent with $s_{2}(t)=\alpha s_{1}(t)$ due to multipath propagation where $\alpha\in\mathbb{C}$. Although the system model in (1) still holds, the pair collapses to a single second-order mode: the source vector becomes $\tilde{\mathbf{s}}(t)=[s_{1}(t),s_{3}(t),\ldots,s_{K}(t)]^{\mathsf{T}}$ and the array response matrix becomes $\tilde{\mathbf{A}}_{\bm{\Omega}}(\bm{\theta})=[\mathbf{a}_{\bm{\Omega}}(\theta_{1})+\alpha\mathbf{a}_{\bm{\Omega}}(\theta_{2}),\mathbf{a}_{\bm{\Omega}}(\theta_{3}),\ldots,\mathbf{a}_{\bm{\Omega}}(\theta_{K})]$, so the coherent pair $\{s_{1}(t),s_{2}(t)\}$ appears as a single source with “combined” steering vector $\mathbf{a}_{\bm{\Omega}}(\theta_{1})+\alpha\mathbf{a}_{\bm{\Omega}}(\theta_{2})$ that generally lies outside the original array response matrix $\mathbf{A}_{\bm{\Omega}}(\bm{\theta})$. As a result, SCM-based MUSIC fails to resolve the pair. By contrast, higher-order statistics (e.g., fourth-order) remain informative under coherence sources because they include per-source self-products such as $\mathbf{a}_{\bm{\Omega}}(\theta_{1})\otimes\mathbf{a}_{\bm{\Omega}}(\theta_{1})$ and $\mathbf{a}_{\bm{\Omega}}(\theta_{2})\otimes\mathbf{a}_{\bm{\Omega}}(\theta_{2})$ that do not collapse, thereby restoring separability [36]. Our proposed snap-TF implicitly leverages such higher-order statistics from non-Gaussian modulated data payloads and remains robust to multipath propagation. Actually, tackling coherent sources in multipath was among the main motivations for introducing higher-order array signal processing [36].
Figure 11: MSE as a function of SNR under coherent sources with $M=5$, $N=9$ and 16QAM symbols.
We present the simulation results in Fig. 11. Specifically, Fig. 11(a) shows the case of $K=3$ sources, where 2 of them are coherent, while Fig. 11(b) simulates the case of $K=9$ sources, where 4 of them are coherent. All SCM-based baselines exhibit an early performance saturation as SNR increases over -5 dB as a consequence of coherent sources. In addition, their performance notably degrades as compared to that under independent sources shown in Fig. 5. By contrast, the proposed Snap-TF algorithm shows an increasingly wide performance advantage at high SNRs, and exhibits a similar performance to that under independent sources, verifying its robustness to multipath propagation.
@@ VI Further Application: Sensing-Assisted MU-MIMO Beam Management
Beam sweeping is an important step of beam management to establish initial alignment [65]. In the downlink, the BS will transmit synchronization signal blocks sequentially across various spatial directions, effectively “sweeping” its coverage range. The users then feedback the index of the strongest beams they receive to the BS to facilitate beam alignment. In the beam sweeping stage, the BS needs to sweep its whole coverage range based on a pre-defined codebook, for which a widely used choice is the discrete Fourier transform (DFT) codebook. Advanced codebook design has also been extensively studied to reduce the beam sweeping overhead and improve the performance.
@@ VI-B Performance
Fig. 13 illustrates the system throughput as a function of the channel coherence time $T_{\text{c}}$ for three beam sweeping strategies. In Fig. 13 (a) and (b), DOA estimation from unknown uplink data symbols is performed at SNRs of -5 dB and 5 dB, respectively. Each data point in the plot is averaged over 10,000 Monte Carlo trials. The proposed codebook + sensing (pruned sweeping) approach consistently outperforms the conventional codebook-only method across all coherence time lengths, with the most pronounced advantage in the short coherence regime. This gain is attributed to the significant reduction of beam sweeping candidates enabled by the sensed uplink DOAs, which effectively lowers training overhead. The beam sweeping overhead for full sweeping is $T_{\text{train}}=64$, while the overhead of sensing pruned sweeping is significantly smaller, only leading to an average sweeping overhead of $T_{\text{train}}=5.0$ in (a) and $T_{\text{train}}=2.9$ in (b), which are respectively $12.8\times$ and $22.1\times$ times fewer compared to full sweeping. As the coherence interval $T_{\text{c}}$ increases, the performance of the full sweeping baseline gradually converges toward that of the proposed sensing-aided pruned sweeping scheme.
@@
VI-C Robustness to Reciprocity Mismatch
In the preceding results we assumed angular reciprocity. In practice, depending on the carrier frequency and propagation environment, slight reciprocity mismatches between uplink and downlink DOAs may occur. In such cases, the proposed sensing-assisted beam management method, specifically the codebook + sensing (pruned sweeping) scheme, remains applicable by appropriately enlarging the width $\delta$ of the search window so that the reciprocity mismatch can be absorbed, thereby preserving the robustness.
To validate this, we conduct a sensitivity test in which the downlink DOA $\theta_{\mathrm{DL},k}$ is modeled the sum of the uplink DOA $\theta_{\mathrm{UL},k}$ and a reciprocity mismatch term $\mathrm{mis}$, given by $\theta_{\mathrm{DL},k}=\theta_{\mathrm{UL},k}+\mathrm{mis}$, where $\mathrm{mis}\sim\mathcal{N}(0,\sigma_{\mathrm{mis}}^{2})$ is modeled as a Gaussian random variable with variance $\sigma_{\mathrm{mis}}^{2}$. We adopt the same simulation setting as Fig. 13(b), except that we now fix the coherence interval as $T_{\text{c}}=200$ symbols, and sweep the strength of reciprocity mismatch between $\sigma_{\mathrm{mis}}\in[0^{\circ},8^{\circ}]$. We compare two different search window widths $\delta\in\{5^{\circ},10^{\circ}\}$.
Figure 14: Robustness of the proposed sensing-assisted beam management to the reciprocity mismatch of uplink and downlink DOAs.
In Fig. 14, we plot the system throughput as a function of $\sigma_{\mathrm{mis}}$. The results show that a larger window, $\delta=10^{\circ}$, makes the pruned sweeping scheme highly robust to reciprocity mismatch. The system throughput remains essentially unchanged for $\sigma_{\mathrm{mis}}\approx 0^{\circ}\thicksim 5^{\circ}$, with only mild degradation beyond this range. In contrast, narrowing the window to $\delta=5^{\circ}$ reduces robustness. The performance begins to decline once $\sigma_{\mathrm{mis}}$ exceeds approximately $2^{\circ}\thicksim 3^{\circ}$. These results suggest that by appropriately setting the search window width $\delta$, sensing-assisted beam management can robustly handle the reciprocity mismatch and outperform existing full sweeping schemes.
VII Conclusion and Future Work
This paper demonstrates the novel paradigm of “sensing for free”—enabling multi-source localization at the BS based on unknown uplink data symbols, without dedicated pilots or additional overhead. Sparse arrays prove particularly promising for this paradigm, as they have the potential to localize more sources than antennas by exploiting their enlarged virtual array. Our attention-only transformer directly processes raw signal snapshots to exploit higher-order statistics for grid-less end-to-end multi-source DOA estimation, circumventing the computational bottlenecks and performance degradation inherent in existing DL-based subspace methods. These accurate DOA estimates can enable significant pruning of downlink beam sweeping candidates, and hence substantially reduce beam training overhead and enhance system throughput at no extra cost. By leaving frame structures and waveforms untouched, the paradigm integrates seamlessly into existing wireless systems and provides inspiration for next-generation 3GPP standardization efforts. Overall, this work reveals how uplink data symbols can be harnessed for free to unlock the substantial sensing capabilities, motivating future research into the untapped potential of regular data symbols for sensing.
Looking ahead, responsible deployment of “sensing for free” requires addressing two important issues: resilience to potential attacks (e.g., location spoofing or intentional interference) and protection of the sensitive positional information. Developing rigorous threat models, validating standard-compliant defense strategies, and establishing privacy-preserving practices and policies in realistic deployments are important research directions. In addition, generalizing the proposed algorithm to wideband and near-field systems is an ongoing future work.